# Multi-Task Label Discovery via Hierarchical Task Tokens for Partially Annotated Dense Predictions

### v1

@@ Appendix C.1 More Quantitative Results
Analysis of Multi-Scale Global Task Token Learning. As we illustrated in Sec. Appendix A.2, we adopt multi-scale global task learning with ResNet-18 backbone on PASCAL-Context. To validate the effectiveness of learning global task tokens on multi-scale features, we compare the performance of Global Task Token Learning with single-scale and multi-scale backbone features respectively in Table 9. As shown in the table, HiTTs with single-scale (SS) Global Token Learning surpass the MTL baseline on both one-label and random-labels settings, with overall $+0.60\%~{}\Delta_{MTL}$ and $+0.97\%~{}\Delta_{MTL}$ on all tasks respectively, and the multi-scale (MS) Global Token Learning further enhances the performance to $+2.84\%~{}\Delta_{MTL}$ and $+3.33\%~{}\Delta_{MTL}$ on all tasks, which indicates the effectiveness of applying global token learning on multi-scale features.
Appendix C.2 More Qualitative Results
@@
Visualization of token distributions. We also provide visualization analysis to show the distributions of fine-grained task tokens on the two tasks of Cityscapes. In Fig. 13, with the aid of OE, the tokens are more orthogonal with each other as shown in the affinity maps, and have better clusters in the feature space as shown in the PCA distributions.
@@
Qualitative prediction comparisons with SOTA works. We additionally provide comparisons with SOTA works on NYUD-v2 and Cityscapes. As shown in Fig. 15, we compare with MTL baseline, XTC li2022learning on NYUD-v2 three tasks, and with MTL baseline, MTAN liu2019end , XTC li2022learning on Cityscapes two tasks. Our method shows clearly better performance in semantic understanding and accurate geometry estimations (including depth and normal estimation), indicating the effectiveness of our method.
@@
Figure 16: Quantitative analysis of the quality of pseudo labels generated by gloabl task tokens.

### v2

@@ B.5.2. Model Analysis
Analysis of Multi-Scale Global Task Token Learning. As we illustrated in Sec. A.4, we adopt multi-scale global task learning with ResNet-18 backbone on PASCAL-Context. To validate the effectiveness of learning global task tokens on multi-scale features, we compare the performance of Global Task Token Learning with single-scale and multi-scale backbone features respectively in Table 12. As shown in the table, HiTTs with single-scale (SS) Global Token Learning surpass the MTL baseline on both one-label and random-labels settings, with overall $+0.60\%~\Delta_{MTL}$ and $+0.97\%~\Delta_{MTL}$ on all tasks respectively, and the multi-scale (MS) Global Token Learning further enhances the performance to $+2.84\%~\Delta_{MTL}$ and $+3.33\%~\Delta_{MTL}$ on all tasks, which indicates the effectiveness of applying global token learning on multi-scale features.
Analysis of Threshold Hyperparameter $\tau_{i}$. We discuss the effect of the confidence threshold hyperparameter, $\tau_{i}$. This threshold is chosen to ensure that high-confidence pixel predictions are masked out to serve as pseudo-labels. To analyze the sensitivity of our method to this hyperparameter, we conduct an ablation study on the NYUD-v2 dataset under the one-label setting. As shown in Table 13, we vary $\tau_{i}$ across a wide range from 0.30.30.3 to 0.950.950.95. The results demonstrate that the performance on all tasks remains remarkably stable with only minor fluctuations. Furthermore, the overall multi-task learning gain ($\Delta_{MTL}$) is consistently and significantly positive across all tested values. This demonstrates that our method is not sensitive to the choice of $\tau_{i}$, showcasing its robustness.
[…]
Implementations with Different Backbones. To demonstrate that our method can be flexibly implemented on different image backbones, we perform additional experiments implementing our HiTTs with a Vision Transformer (ViT (Dosovitskiy et al., 2020)) backbone. The results on NYUD-v2 are presented in Table 16. The quantitative results clearly show that when equipped with the more powerful ViT-base backbone, our method achieves a substantial performance improvement across all tasks under both the one-label and random-labels settings. For instance, in the random-labels setting, using ViT-base boosts the Semantic Segmentation mIoU from 47.73 to 61.88. This not only confirms the flexibility of our approach but also highlights its potential to achieve even greater performance when paired with more advanced backbone architectures.
Analysis of Incorporating Dense FixMatch (Pieropan et al., 2022) on NYUD-v2. Our method can also incorporate general semi-supervised dense prediction strategies, e.g. Dense FixMatch (Pieropan et al., 2022). The quantitative results on the NYUD-v2 one-label setting are shown in Table 14. The results indicate that Dense FixMatch provides consistent performance improvements. This demonstrates the efficacy of Dense FixMatch as a versatile component for enhancing various dense prediction tasks in a semi-supervised context.
Table 13. Ablation study on the hyperparameter $\tau_{i}$.
$\tau_{i}$ Semseg. Depth. Normal. $\Delta_{MTL}$
mIoU mErr mErr (%)↑
0.95 46.67 0.4585 25.48 3.11
0.90 47.10 0.4573 25.43 3.57
0.70 46.89 0.4621 25.36 3.17
0.50 47.00 0.4690 25.35 2.79
0.30 47.11 0.4706 25.33 2.78
Table 14. Comparisons with incorporating general semi-supervised dense prediction method (Pieropan et al., 2022) on NYUD-v2 one-label setting.
   Method    Semseg.    Depth.    Normal.
   mIoU    mErr    mErr
   Ours    47.30    0.4539    25.40
   Ours+(Pieropan et al., 2022)    47.72    0.4502    25.33
Table 15. Comparison on NYUD-v2 under the settings of DiffusionMTL (Ye and Xu, 2024). (P) and (F) represent the Prediction Diffusion and Feature Diffusion modes for DiffusionMTL (Ye and Xu, 2024). “*” denotes the re-implemented results to align the settings. Our method performance outperforms previous methods.
Setting Model Semseg. Depth. Normal. $\Delta_{m}$
mIoU $\uparrow$ absErr $\downarrow$ mErr $\downarrow$ (%) $\uparrow$
One-Label
STL 45.28 0.4802 25.93 -
MTL baseline 43.92 0.5138 26.44 -3.99
SS (Li et al., 2022) 27.52 0.6499 33.58 -
XTC (Li et al., 2022) 30.36 0.6088 32.08 -
XTC* (Li et al., 2022) 43.97 0.5140 26.30 -3.79
DiffusionMTL (P) (Ye and Xu, 2024) 44.97 0.5137 26.17 -2.86
DiffusionMTL (F) (Ye and Xu, 2024) 44.47 0.5059 25.84 -2.27
Ours* 47.30 0.4539 25.40 3.99
Random-Labels
STL 48.25 0.4792 24.65 -
MTL baseline 45.93 0.4839 25.53 -3.12
SS (Li et al., 2022) 29.50 0.6224 33.31 -
XTC (Li et al., 2022) 34.26 0.5787 31.06 -
XTC* (Li et al., 2022) 46.03 0.4811 25.97 -3.44
DiffusionMTL (P) (Ye and Xu, 2024) 47.44 0.4803 25.26 -1.45
DiffusionMTL (F) (Ye and Xu, 2024) 46.82 0.4743 24.75 -0.77
Ours* 47.73 0.4510 24.86 1.32
Table 16. Ablation study of our method with different backbones on NYUD-v2. The results show that our method is flexible and achieves significant gains when paired with a more powerful ViT backbone.
Setting Backbone Semseg. Depth. Normal.
mIoU $\uparrow$ absErr $\downarrow$ mErr $\downarrow$
One-Label ResNet-18 47.30 0.4539 25.40
ViT-base 58.38 0.3740 23.65
Random-Labels ResNet-18 47.73 0.4510 24.86
ViT-base 61.88 0.3979 23.03
Figure 10. Comparison of the task confidence map before and after refined by the fine-grained task tokens, which greatly encourage high-confidence predictions on all tasks (red color represents high-confidence areas). For noisy data like the second photo taken in a dark environment, this enhancement are more significant.
B.6. More Qualitative Results
@@
Role of Fine-grained Task Tokens. To illustrate the role of our fine-grained task tokens, we visualize task confidence maps before and after the refinement process in Fig. 10. The fine-grained tokens significantly enhance prediction confidence across all tasks, as indicated by the expansion of high-confidence areas (represented in red). This enhancement is particularly evident in challenging scenarios, such as the noisy image captured in a dark environment (second row), where the refined map shows a marked improvement in clarity and confidence.
Visualization of token distributions. We also provide visualization analysis to show the distributions of fine-grained task tokens on Cityscapes. In Fig. 11, with the aid of OE, the self-correlation map of tokens will be more diagonal, and the distributions after PCA have better clusters in 3-dimensional feature space.
@@
Qualitative prediction comparisons with SOTA works. We additionally provide comparisons with SOTA works on NYUD-v2 and Cityscapes. As shown in Fig. 13, we compare with MTL baseline, XTC (Li et al., 2022) on NYUD-v2 three tasks, and with MTL baseline, MTAN (Liu et al., 2019), XTC (Li et al., 2022) on Cityscapes two tasks. Our method shows clearly better performance in semantic understanding and accurate geometry estimations (including depth and normal estimation), indicating the effectiveness of our method.
@@
Figure 11. Visualization of self-affinities heatmap (left) and PCA for distributions (right) of fine-grained task tokens of the two tasks on NYUDv2 and Cityscapes validation sets. With orthogonal embeddings, the affinities between different tokens are low and the clustering of token distributions on each category is better.
Figure 12. Comparisons of task score maps produced by global task tokens and fine-grained task tokens. The upper part is the visualization of samples on NYUD-v2 while the lower part is on Cityscapes.
Figure 13. Comparisons with SOTA works on NYUD-v2 (upper part) and Cityscapes (lower part).
Figure 14. Quantitative analysis of the quality of pseudo labels generated by gloabl task tokens.